OpenStudy (anonymous):

how can i compare 12 9 -- and -- 16 12

OpenStudy (anonymous):

\[\frac{12}{16}=\frac{3}{4}\] and \[\frac{9}{12}=\frac{3}{4}\]

OpenStudy (angela210793):